# 1]FAIR at Meta

### v1

@@ 1 Introduction
Over the past few years, Transformer-based architectures (Vaswani et al., 2017) have demonstrated impressive performance in different tasks, including holding conversations at the human level (Shuster et al., 2022; OpenAI, 2022, 2023; Touvron et al., 2023), high-quality image understanding (Caron et al., 2021; Oquab et al., 2024; Assran et al., 2023) and video generation (Singer et al., 2023), multi-modal generation (Girdhar et al., 2023; Radford et al., 2021), and code completion (Roziere et al., 2023; OpenAI, 2021). By training these architectures with a huge amount of data, the resulting models, such as Large Language Models (LLMs), can generalize well in real-world use cases.
Despite these successes, Transformer-based architectures and LLMs still struggle when it comes to solving planning and reasoning tasks. Previous studies (Momennejad et al., 2023; Valmeekam et al., 2023a, b) demonstrate that LLMs fall short in multi-step planning tasks (Valmeekam et al., 2023b) or when performing higher-order reasoning (Momennejad et al., 2023; Fan et al., 2020).
In recent years, various methods have been proposed to improve the performance of Transformers in reasoning and planning tasks. One common and effective approach is to simulate the human thinking process and produce intermediate “thoughts” before outputting a response. Chain-of-Thought (CoT) prompting (Wei et al., 2022) encourages the model to predict the intermediate steps and to “think” step by step. Tree-of-thoughts (ToT) uses a branching strategy and critics to generate different thought paths before picking the best one (Yao et al., 2023). While these techniques are often effective, there are studies showing that in many cases, they may lead to worse performance, for example due to self-enforcing (Huang et al., 2023). Furthermore, techniques effective on one dataset may not work well on others due to changes in the type of reasoning involved (e.g., spatial reasoning vs. mathematical reasoning vs. common-sense reasoning). How to enable Transformers or LLMs to plan, solve multi-step decision making tasks, and perform different types of reasoning still remains elusive and an active area of research.
These methods stand in sharp contrast with traditional symbolic planning and search techniques. While such techniques may not exhibit the language understanding capabilities of LLMs trained on internet-scale datasets, they are well-defined and can solve planning tasks that are far more complex than the tasks state-of-the-art LLMs can solve (Momennejad et al., 2023; Valmeekam et al., 2023a). Additionally, there often exist formal guarantees for solutions computed by these traditional methods, because symbolic planning algorithms often follow a well-defined rule-based search procedure.
@@
In this paper, we demonstrate how to train Transformers to solve complex planning tasks and present Searchformer, a Transformer model that computes an optimal plan in fewer search steps than symbolic planning algorithms such as $A^{*}$ search (Russell and Norvig, 2021, Chapter 3) in multi-step planning tasks such as maze navigation and solving Sokoban puzzles. To achieve this, we introduce search dynamics bootstrapping, a method that first trains a Transformer model to imitate the search procedure of $A^{*}$ and then fine-tunes the model to find an optimal plan with fewer search steps.
In the first step a Transformer model is trained to imitate $A^{*}$ search. Here, we generate synthetic datasets by running $A^{*}$ search in randomly generated planning task instances. While $A^{*}$ is executed, we log the performed computation and optimal plan into a sequence of words, called tokens. The resulting training dataset then contains an execution trace of $A^{*}$ and encodes information about the search dynamics of $A^{*}$ itself. A Transformer model is then trained to generate these token sequences along with an optimal plan for any given planning task.
In the second step, we show that our Searchformer trained with search-augmented sequences—sequences that include $A^{*}$’s execution trace—can be further improved via expert iteration (Gulcehre et al., 2023), a method that encourages the Transformer to generate fewer search steps while still generating optimal solutions. This procedure leads to a neural planning algorithm that is implicitly encoded in the Transformer’s network weights, and finds an optimal plan with high probability in fewer search steps than standard $A^{*}$ search. Specifically, when solving Sokoban puzzles, our models solve 93.7% of all test tasks while performing search steps that are on average 26.8% shorter than $A^{*}$ search. This paves the way to use Transformers to move beyond traditional symbolic planning algorithms.
To better understand how the used training data and number of model parameters influence performance of the resulting models, we present a set of ablation studies. We train models on both solution-only sequences (token sequences including task description and final plan) and search-augmented sequences (token sequences including task description, search tree dynamics, and final plan) and compare their performance. Despite the fact that search-augmented sequences are thousands of tokens long, search-augmented models (that are trained on search-augmented sequences) can still learn to generate the correct search dynamics and optimal plans with high probability. Further, compared to solution-only models (that are trained on solution-only sequences), search-augmented models generate an optimal plan more often on unseen tasks with ten times fewer training sequences, highlighting the power of including $A^{*}$’s search dynamics into the training process of Transformer models. Since the search dynamics can be included into the training data simply by through logging when generating synthetic training sequences, our study provides a data-efficient way to improve the reasoning power of Transformer models. We also find that solution-only models do not improve with more parameters, showing that the type of training data used plays a critical role in this setting.
@@ 2 Related Work
While existing work (Trinh et al., 2024; Ruoss et al., 2024) leverages synthetic datasets to learn strong policies for reasoning, our study is fundamentally different in this regard. We focus on improving the reasoning capability embedded in a Transformer’s weights, while algorithms such as AlphaZero (Silver et al., 2018), MuZero (Schrittwieser et al., 2020) and AlphaGeometry (Trinh et al., 2024) treat general neural network models (including Transformers) as black-boxes, and use existing symbolic planning techniques to improve the neural network itself. For example, Silver et al. (2017) use MCTS as a policy improvement operator to update the neural network’s weights. In contrast, the search dynamics bootstrapping method in this work uses a Transformer model to generalize towards more efficient search patterns and improve the model itself. A planning algorithm is only used to initially train a Transformer model.
ScratchPad (Nye et al., 2021) train neural networks on execution traces of reasoning tasks, Yang et al. (2022), Pallagani et al. (2022), and Ruoss et al. (2024) train models to predict the optimal solution or an optimal action directly. In contrast, we find more efficient LLM-based planning algorithms by predicting the search dynamics of a planning algorithm, which is the execution trace when the solver attempts to find the optimal solution, which is a harder problem than predicting an optimal decision making strategy.
Our work also bears some similarity with neuro-symbolic systems (e.g., Neural Turing Machines (Graves et al., 2014), Neural Programming Architecture (Cai et al., 2017)), which build differentiable architectures to mimic the functionality of existing symbolic systems. However, these methods use dedicated components (e.g., explicit memory components, built-in recursion), while Searchformer uses a general Transformer architecture. By using long contexts and positional embeddings (Chen et al., 2023; Peng et al., 2023) to predict an optimal plan, Searchformer performs well in solving different complex planning tasks. Ultimately, our work is a step towards building more general neural architectures that automatically learn a reasoning and planning mechanism without prior human knowledge.
@@
Promptbosstart 0 2goal 1 0wall 1 2wall 2 0eosResponsebosplan 0 2plan 0 1plan 1 1plan 1 0eos
@@
[…]
@@
Figure 1: Expressing a planning task in token sequences. 1(a): A $3\times 3$ maze navigation task where an agent must navigate from the start cell to the goal cell without entering a wall cell. For each transition, a cost of one is given and the optimal plan is the shortest path from start to goal. 1(b): The $3\times 3$ maze navigation task is expressed as a prompt token sequence (left panel) and the optimal plan is expressed as a response token sequence (right panel). The start and end of a response sequence is indicated by a beginning-of-sequence token, bos, and an end-of-sequence tokens, eos. The numbers indicate $x,y$ coordinates. 1(c): The search dynamics of the $A^{*}$ algorithm (left panel) is logged into an execution trace represented by a token sequence (right panel). For maze tasks, coordinates are encoded with integer tokens. The next two numbers costs associated with that node (letter “c” is added to differentiate them from coordinate numbers). Each token is mapped to a unique non-negative integer and this integer is processed in the neural network.
Using Transformer architectures to solve complex sequential decision making has been studied in prior work in a reinforcement learning (RL) setting (Chen et al., 2021; Janner et al., 2021; Laskin et al., 2023). Decision Transformer (Chen et al., 2021) predicts the next optimal action given a trajectory segment that is annotated with total return values, while Trajectory Transformer (Janner et al., 2021) focuses on learning a world model that predicts future states, actions, and rewards. Similarly, in Algorithm Distillation (Laskin et al., 2023) a Transformer architecture is trained to predict the next action given a trajectory segment of trial and error interactions as context. In contrast, we incorporate how a planning algorithm arrives at a final result into a Transformer models. MCTSNet (Guez et al., 2018) also attempts to learn the search procedure itself, but still hard-codes the MCTS search procedure into neural network, which leads to quadratic backpropagation overhead and can only deal witih up to 500 rollouts, while our approach can deal with much longer search execution trace. We demonstrate that Transformers can not only imitate a symbolic planning algorithm but also improve past the algorithm it was initially trained on via fine-tuning.
@@ 3 Problem Setup
Figure 2: Example Sokoban puzzle [^2] This level image was composed using image icons from https://github.com/morenod/sokoban (accessed 2023-11-21).. The prompt, $A^{*}$ execution trace, and optimal plan for this task is illustrated in Figure 10 in Appendix A.3. For Sokoban, we only use a non-deterministic $A^{*}$ implementation.
We consider two domains: maze navigation tasks (Figures 1(a)) and solving Sokoban puzzles (Figure 2). In maze navigation, the goal is to find the shortest path in an $n$-by-$n$ maze between a start location and a goal location without passing through wall cells. Figure 2 depicts a Sokoban puzzle where the worker can move up, down, left, or right to push each box onto a dock (crossed grid cells). The worker cannot pull a box away from a wall cell. To solve the puzzle, each box must be moved onto a dock in as few steps as possible. An incorrect move may immediately lead to a dead end and thus careful planning is needed to solve the task. Furthermore, each state in a Sokoban puzzle consists of every possible combination of a box and worker position. Consequently, state spaces are large making a Sokoban puzzle computationally more difficult to solve for algorithms such as $A^{*}$ search.
@@
  • A frontier set containing the current search frontiers;
@@
In the maze example in Figure 1(a), each node corresponds to an empty (non-wall) grid cell. For each node, the algorithm computes a heuristic value and a cost from start value. At any given iteration, which node is searched next is determined by the content of the frontier and closed sets as well as the heuristic and cost from start values (Figure 1(c), left panel). The $A^{*}$’s execution trace is collected by tracking all insertion operations into the frontier and closed set along with heuristic and cost from start values. The right panel in Figure 1(c) illustrates the resulting trace for the maze example shown in Figure 1(b). Each row corresponds either to an insertion of a node into the frontier, indicated by a create token, or to moving a node into the closed set, indicated by a close token. Each node is represented by its $(x,y)$ position in the maze as well as the two cost tokens. The resulting plan is then appended to this trace. This trace is constructed such that given any prefix of this trace the next token can be predicted correctly. For the maze datasets, $A^{*}$ uses the Manhattan distance to the goal location as a heuristic. In Sokoban, $A^{*}$ first matches every box to the closest dock and then computes the sum of all Manhattan distances between each box and dock pair.
@@
  • Search-augmented sequence. This variant contains sequences of the format <prompt><trace><plan>, where the <trace> part encodes $A^{*}$ execution trace and <plan> encodes the optimal plan.
  • Solution-only sequence. This variant omits $A^{*}$’s exeuction trace and only contains sequences of the format <prompt><plan>.
For example, the $3\times 3$ maze navigation task (Figure 1(a)) can be mapped into a prompt and plan sequence (Figure 1(b)). The <trace> part is constructed by running $A^{*}$ on each planning task and logging the performed computations into a token sequence (Figure 1(c)). While the optimal plan in both cases can be fairly short, the corresponding execution traces can be quite long, in particular when $A^{*}$’s heuristic underestimates the cost to reach the goal state and often misleads the search to explore in-feasible paths.
As illustrated in Figure 1, all token sequences are synthetically generated and use a special vocabulary. Furthermore, every model is trained from scratch—no natural language dataset is used for training. Consequently, the resulting models are specifically trained to only predict sequences that outline optimal plans for a set of different planning tasks. After training, the model’s output is parsed and evaluated if it contains a correct and optimal solution plan.
@@ 3.2 Training a Transformer model
To train a Transformer model, we randomly generate a set of planning tasks together with their corresponding token sequences. In each dataset, each task is unique and the test set is constructed such that it does not contain any duplicate of the training set. For example, a maze task that occurs in the training dataset does not occur in the test dataset and no task is repeated. With this experiment design, we hope to gain insight into how Transformers can be used to solve planning tasks.
By including intermediate computation steps, the Transformer model is trained to effectively imitate the computation performed by the $A^{*}$ algorithm. This approach to training a neural network is similar to Procedure Cloning (Yang et al., 2022) where a neural network is not only presented with input-output pairs (in our case task prompts and optimal plans), but also with examples of how to “think” about a particular task
In our experiments, we train an adaptation of the encoder-decoder T5 architecture (Raffel et al., 2020) and use Rotary Position Embeddings (RoPE) (Su et al., 2023). More details and hyper-parameters about the models we use can be found in Appendix A.2. The encoder processes the <prompt> part of a training sequence, and the decoder processes either a <trace><plan> formatted sequence to train a search-augmented model or only a <plan> formatted sequence when training a solution-only model. Depending on the model variant, each network is trained to maximize the cross-entropy between the distribution of the decoder generations and the distribution of sampling a corresponding sequence from the training dataset. Gradient updates are computed using teacher forcing. Appendix A.1 describes our optimization setup in more detail.

### v2

@@ Abstract
While Transformers have enabled tremendous progress in various application settings, such architectures still trail behind traditional symbolic planners for solving complex decision making tasks. In this work, we demonstrate how to train Transformers to solve complex planning tasks. This is accomplished by training an encoder-decoder Transformer model to predict the search dynamics of the $A^{*}$ search algorithm. We fine tune this model to obtain a Searchformer, a Transformer model that optimally solves previously unseen Sokoban puzzles 93.7% of the time, while using up to 26.8% fewer search steps than the $A^{*}$ implementation that was used for training initially. In our training method, $A^{*}$’s search dynamics are expressed as a token sequence outlining when task states are added and removed into the search tree during symbolic planning. Searchformer significantly outperforms baselines that predict the optimal plan directly with a 5–10$\times$ smaller model size and a 10$\times$ smaller training dataset. Lastly, we demonstrate how Searchformer scales to larger and more complex decision making tasks with improved percentage of solved tasks and shortened search dynamics.
@@ 1 Introduction
Transformer-based architectures (Vaswani et al., 2017) have demonstrated impressive performance in different tasks, including holding conversations at the human level (Shuster et al., 2022; OpenAI, 2022, 2023; Touvron et al., 2023), high-quality image understanding (Caron et al., 2021; Oquab et al., 2024; Assran et al., 2023) and video generation (Singer et al., 2023), multi-modal generation (Girdhar et al., 2023; Radford et al., 2021), and code completion (Roziere et al., 2023; OpenAI, 2021). By training these architectures on internet-scale datasets, the resulting models, such as Large Language Models (LLMs), can generalize well in real-world use cases.
Despite these successes, Transformer-based architectures and LLMs still struggle when it comes to solving planning and reasoning tasks. Previous studies demonstrate that LLMs fall short in multi-step planning tasks (Valmeekam et al., 2023a, b) or when performing higher-order reasoning (Momennejad et al., 2023; Fan et al., 2020).
In recent years, various methods have been proposed to improve the performance of Transformers in these settings. One approach is to simulate the human thinking process and produce intermediate “thoughts” before outputting a response. Chain-of-Thought (CoT) prompting (Wei et al., 2022) and the Tree-of-thoughts (ToT) method (Yao et al., 2023) encourage the model to “think” step by step. While these techniques are often effective, they can also lead to worse performance, for example due to self-enforcing (Huang et al., 2023). Furthermore, techniques effective on one dataset may not work well on others due to changes in the type of reasoning involved (e.g., spatial reasoning vs. mathematical reasoning). How to enable Transformers and LLMs to plan, solve multi-step decision making tasks, and perform reasoning still remains elusive and an active area of research.
@@
We demonstrate how to train Transformers to robustly solve complex planning tasks. Similar to LLMs, we train Transformers to predict the next word given a sequence of words. Our experiments use synthetically generated datasets with a synthetic language and vocabulary. Using this framework, we demonstrate how to construct training data such that the resulting model imitates the computation performed by $A^{*}$ search (Russell & Norvig, 2021, Chapter 3). Lastly, we present Searchformer, a Transformer model that solves complex planning tasks in fewer search steps than our $A^{*}$ reference implementation. This model is obtained through search dynamics bootstrapping, a method that first trains a Transformer to imitate $A^{*}$’s search process and then fine-tunes the model to find an optimal plan within fewer search steps.
To train a Transformer to perform planning, we express a planning task and its optimal solution plan as a sequence of words, called tokens. We also log the computation performed by $A^{*}$ into an execution trace consisting of a token sequence, resulting in a sequence dataset that captures $A^{*}$’s search dynamics. Using these search-augmented sequences, a Transformer model is trained to generate token sequences that encode $A^{*}$’s search dynamics along with an optimal plan.
Subsequently, the resulting search-augmented model is fine-tuned to generate shorter token sequences while still outputting an optimal plan. We refer to this final fine-tuned model as a Searchformer. When solving Sokoban puzzles, our models solve 93.7% of all test tasks while performing on average 26.8% fewer search steps than our $A^{*}$ reference implementation.
Through a sequence of experiments that control task complexity, dataset size, and model size, we demonstrate that including execution traces into the training data increases performance on an independent test task set—despite a 10×10$\times$10 × to 100×100$\times$100 × increase in the length of the generated sequences. We find that search-augmented models (that include execution traces into their training data) generate an optimal plan more often on unseen tasks with ten times fewer training sequences than a larger solution-only model (that is trained on sequences only including a task description and task solution). This result highlights the power of including $A^{*}$’s search dynamics into the training process of Transformer models.
@@ 2 Related Work
While existing work (Trinh et al., 2024; Ruoss et al., 2024) leverages synthetic datasets to learn policies for reasoning, our study is fundamentally different in this regard. We focus on improving the reasoning capability embedded in a Transformer’s weights. Existing algorithms such as AlphaZero (Silver et al., 2018), MuZero (Schrittwieser et al., 2020), and AlphaGeometry (Trinh et al., 2024) optimize a neural network using the output of existing symbolic planning algorithms, whose internal states are not used (i.e., treated as black-boxes). For example, Silver et al. (2017) use MCTS as a policy improvement operator to update the neural network’s weights. In contrast, the presented search dynamics bootstrapping method uses a Transformer model to generalize towards more efficient search patterns and improves the model itself. A planning algorithm (together with its internal search dynamics) is used to initially train a Transformer model.
Prior work focuses on training a neural network on execution traces of reasoning tasks (Nye et al., 2021) or on training a neural network to predict an optimal action (Yang et al., 2022; Pallagani et al., 2022; Ruoss et al., 2024). In contrast, we focus on training a Transformer to generate the entire search process of $A^{*}$ when computing an optimal plan. Instead of only predicting a single action, our model predicts an entire multi-step plan to solve a task.
Our work bears some similarity with neuro-symbolic systems (Graves et al., 2014; Cai et al., 2017), which build differentiable architectures to mimic the functionality of existing symbolic systems. However, these methods use dedicated components (e.g., explicit memory components, built-in recursion), while Searchformer focuses on next-token prediction. Here, Searchformer relies on generating long contexts and position embeddings (Chen et al., 2023; Peng et al., 2023) to predict in optimal plan. Ultimately, our work sheds light on how to build more general architectures that automatically learn a planning mechanism.
Using Transformer architectures to solve complex sequential decision making tasks has been studied in prior work in a reinforcement learning (RL) setting (Chen et al., 2021; Janner et al., 2021; Laskin et al., 2023). However, this prior work presents different methods to modelling trajectories of trial and error interactions and focuses on predicting a next action, state, or rewards or a combination of them. In contrast, we demonstrate how to use a Transformer to model the search steps involved in computing an optimal multi-step plan. MCTSNet (Guez et al., 2018) also attempts to learn the search procedure itself, but still hard-codes the MCTS search procedure (Coulom, 2006) into a neural network, which leads to quadratic backpropagation overhead and can only deal with up to 500 step rollouts, while our approach can deal with much longer search execution trace. We demonstrate that Transformers can not only imitate a symbolic planning algorithm but can also be used to discover more efficient heuristics via fine-tuning.
@@
Promptbosstart  0 2goal   1 0wall   1 2wall   2 0eosResponsebosplan  0 2plan  0 1plan  0 0plan  1 0eos
@@
A* planning algorithm 0: Start node $n_{\text{start}}$ and goal node $n_{\text{goal}}$. 1:$\mathcal{S}_{\text{closed}}\leftarrow\{\}$ 2:$\mathcal{S}_{\text{frontier}}\leftarrow\{n_{\text{start}}\}$ 3:while  $|\mathcal{S}_{\text{frontier}}|>0$  do 4:    $n_{\text{curr}}=\arg\min_{n\in\mathcal{S}_{\text{frontier}}}\text{cost}(n)$ 5:    $\mathcal{S}_{\text{closed}}\leftarrow\mathcal{S}_{\text{closed}}\cup\{n_{\text{curr}}\}$ 6:    for  $n_{\text{child}}\in\text{children}(n_{\text{curr}})$  do 7:       if $\text{pos}(n)=\text{pos}(n_{\text{child}})$ for any $n\in\mathcal{S}_{\text{closed}}\cup\mathcal{S}_{\text{frontier}}$ then 8:          if $c(n)+h(n)\leq c(n_{\text{child}})+h(n_{\text{child}})$ then 9:             continue 10:          end if 11:       end if 12:       Set $\text{parent}(n_{\text{child}})\leftarrow n_{\text{curr}}$ 13:       $\mathcal{S}_{\text{frontier}}\leftarrow\mathcal{S}_{\text{frontier}}\cup\{n_{\text{child}}\}$ 14:    end for 15:end while 16: Compute and return plan by recursing on parents of $n_{\text{curr}}$. Tokenization of algorithm executionAdd node to frontierAdd node to closed setCost from start $c(n)$Heuristic value $h(n)$boscreate 0 2 c0 c3close  0 2 c0 c3create 0 1 c1 c2close  0 1 c1 c2create 0 0 c2 c1create 1 1 c2 c1close  0 0 c2 c1create 1 0 c3 c0close  1 0 c3 c0plan    0 2plan   0 1plan   0 0plan   1 0eosTracePlan
@@
Figure 1: Expressing a planning task in token sequences. 1(a): A $3\times 3$ maze navigation task where the goal is to find a the shortest path from start to goal without entering a wall cell. 1(b): The $3\times 3$ maze navigation task is expressed as a prompt token sequence (left panel) and the optimal plan is expressed as a response token sequence (right panel). The start and end of a sequence is indicated by a beginning-of-sequence token, bos, and an end-of-sequence token, eos. Numbers indicate $x,y$ coordinates. 1(c): The search dynamics of the $A^{*}$ algorithm (left panel) is logged into an execution trace (right panel). The last two tokens in the trace encode the cost-since-start value $c(n)$ and the heuristic value $h(n)$ (letter “c” distinguishes costs from coordinate numbers). The $A^{*}$ algorithm is described in detail by Russell & Norvig (2021, Chapter 3).
@@ 3 Problem Setup
Figure 1 provides an overview of our synthetic dataset generation process. We consider two domains: maze navigation (Figure 1(a)) and solving Sokoban puzzles (Figure 5 in Appendix C). In maze navigation, the goal is to find the shortest path through an $n$-by-$n$ maze. In Sokoban, a worker can move up, down, left, or right and has to push each box onto a dock to solve the puzzle. An incorrect move may immediately lead to a dead end and thus reasoning across multiple time steps is required to solve the puzzle. Each state in a puzzle consists of a combination of box and worker positions, making Sokoban computationally more difficult to solve than maze navigation.
@@
  • A frontier set containing the current search frontiers.
@@
In the maze example in Figure 1(a), each node corresponds to an empty (non-wall) grid cell. For each node, the algorithm computes a heuristic value and a cost from start value. At any given iteration, which node is searched next is determined by the content of the frontier and closed sets as well as the heuristic and cost from start values (Figure 1(c), left panel). $A^{*}$’s execution trace is collected by tracking all insertion operations into the frontier and closed set along with heuristic and cost from start values (Figure 1(c), right panel). The right panel in Figure 1(c) illustrates the resulting trace for the maze example shown in Figure 1(b). Each row corresponds either to an insertion of a node into the frontier, indicated by a create token, or to moving a node into the closed set, indicated by a close token. Each node is represented by its $(x,y)$ position in the maze as well as the two cost tokens. The resulting plan is then appended to this trace. This trace is constructed such that given any prefix the next token can be predicted correctly. For the maze datasets, $A^{*}$ uses the Manhattan distance to the goal location as a heuristic. In Sokoban, $A^{*}$ first matches every box to the closest dock and then computes the sum of all Manhattan distances between each box and dock pair.
@@
  • Solution-only sequences of the format <prompt><plan>, where the <prompt> part encodes a task description and the <plan> part the optimal plan (Figure 1(b)).
  • Search-augmented sequences of the format <prompt><trace><plan>, where the <trace> part encodes $A^{*}$’s execution trace (Figure 1(c)).
Because every model is trained from scratch, the resulting models are specifically trained to only predict sequences that outline optimal plans for a set of different planning tasks. After training, the model’s output is parsed and evaluated if it contains an optimal or feasible solution plan.
@@ 3.2 Training a Transformer model
When generating a token sequence dataset, each task is unique and the test set is constructed such that it does not contain any duplicate of the training set. With this experiment design, we hope to gain insight into how Transformers can be used to solve planning tasks and generalize to previously unseen test tasks.
By including intermediate computation steps, the Transformer model is trained to effectively imitate the computation performed by the $A^{*}$ algorithm. Different from Procedure Cloning (Yang et al., 2022) where a neural network is learned to predict the optimal state/action sequence (in our case task prompts and optimal plans), our Transformer model also learns to predict the entire thinking process, including the attempted but failed paths, that leads to the optimal plan.
For each experiment an adaptation of the encoder-decoder T5 architecture (Raffel et al., 2020) is used that integrates Rotary Position Embeddings (RoPE) (Su et al., 2023). More details and hyper-parameters can be found in Appendix B. The encoder processes the <prompt> part of a training sequence, and the decoder processes either a <trace><plan>-formatted sequence (search-augmented model) or only a <plan>-formatted sequence (solution-only model). Depending on the model variant, each network is trained to maximize the cross-entropy between the distribution of the decoder generations and the distribution of sampling a corresponding sequence from the training dataset. Appendix A describes our optimization setup in more detail.